B.Calculate the boiling point and freezing point of the following solutions:

1.0.575 molal aqueous solution

2.10 g of C6 H12 O6 dissolved in 259 g of water

3.0.5 mole Br2 dissolved in 507 g chloroform


D.A solution of propylene glycol (antifreeze) in the car’s radiator has a freezing point of -20 degree celcius.What is the molarity of the solution?


E.The boiling point of an aqueous solution is 102.5 degree celcius. What is the freezing point of the solution?

 


1
Expert's answer
2022-03-07T13:47:01-0500

B.

Taking "k_b" for water "=0.515\u00b0C\/M"

and "k_f" for water "=1.86\u00b0C\/M"


1.

"\Delta\>T_b=0.515×0.575=0.30"

Boiling point "=100+0.3=100.3\u00b0C"


"\Delta\>T_f=1.86×0.575=1.1"

Freezing point "=0-1.1=-1.1\u00b0C"


2.

Molar mass of C6H12O6 "=180g\/mol"


Concentration "=\frac{1000}{259}×\frac{10}{180}=0.2145M"



"\Delta\>T_b=0.2145×1.86=0.40"

Freezing point "=0-0.40=-0.40"